Consider the following balanced chemical reaction:

C + 2 Cl2 → CCl4

What mass of CCl4 (in grams) will be produced when 9.2 g of carbon reacts with 37.8 g of chlorine? HINT: Consider which compound is the limiting reagent.

Express your answer to 2 significant figures.
1
Expert's answer
2019-09-23T03:48:50-0400

C + 2 Cl2 → CCl4

nC= 9.2 / 12 =0.767 moles.

nCl= 37.8/71= 0.532 moles.

Mol ratio = 1:2= 0.767:0.532/2.

Therefore, Cl2 is the limiting reagent.

(0.532/2) x 154 = 40.964 g (CCl4).